# LORE: A Large Generative Model for Search Relevance

### v1

@@ 3.3.3 Distillation by SFT
Fig 9 illustrates the trends in model metrics as the data proportion increases. The results demonstrate that the model can rapidly acquire the correct reasoning format in the early training stage, achieving a format accuracy exceeding 98% with only 20% of the training data. The Pass@8 metric exhibits rapid growth in the early training phase but subsequently displays a phenomenon of diminishing marginal returns. When approximately 60% of the training data is utilized, the growth rate of this metric plateaus, showing no significant improvement. Based on these experimental observations, this study selects the 40% data proportion as the optimal training scale, with which the cold-start model is trained.
@@ 3.4.3 Training Strategy Optimization
Dynamic Sampling.(DAPO) As RL training progresses, a sample effectiveness decay problem emerges: after the model policy converges, an increasing number of samples exhibit consistent outcomes across multiple sampling attempts (either all correct or all incorrect), resulting in an advantage function $A_{i}=0$. Although these saturated samples participate in training, they fail to generate effective gradients, leading to computational resource waste and slowing down convergence speed. To address this issue, we introduce a dynamic sampling strategy: during the Rollout phase, we calculate sample advantage values in real-time and discard zero-advantage samples. A policy update round is only triggered when the accumulated number of non-zero advantage samples reaches the preset batch size.
Entropy Collapse Optimization. During RL training, as the policy gradually converges, the entropy of the model’s output distribution rapidly decreases, causing generation results to become deterministic. This premature convergence phenomenon limits the model’s exploration capability, making it difficult to discover better reasoning paths and ultimately constraining the improvement of the performance upper bound. To address this issue, we explored multiple strategies to slow down the entropy decline and maintain the model’s exploration capability.
@@
(2) On-policy. Traditional off-policy methods perform multi-epoch training on the same batch of experience data, using importance sampling constraints to prevent excessive policy deviation. However, this conservative update mechanism limits the model’s exploration capability. To address this, we adopt a strictly on-policy strategy: each batch of sampled data ($batch_{s}ize\texttimes group_{s}izesamples$) is discarded after being used for only one gradient update, and the next update uses freshly sampled data from the new policy.
@@
$\begin{split}\mathcal{J}_{\text{GRPO}}^{2}(\theta)=&\mathbb{E}_{x\sim\mathcal{D},\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|x)}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|y_{i}|}\sum_{t=1}^{|y_{i}|}\min\left(w_{i,t}(\theta)\widehat{A}_{i,t},\text{clip}(w_{i,t}(\theta),1-\varepsilon,1+\varepsilon)\widehat{A}_{i,t}\right)\right]\\ &\quad-\alpha\cdot\mathbb{E}_{x\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta}(\cdot|x)}\left[-\frac{1}{G}\sum_{i=1}^{G}\log\pi_{\theta}(o_{i}|x)\right]\end{split}$ (20)
@@
Fig 13 presents a comparison of several methods. The Clip-higher strategy exhibits optimal performance by effectively slowing the decline rate of policy entropy while maintaining its decreasing trend, achieving a favorable balance between exploration and exploitation, and ensuring training stability alongside significant performance improvements. The On-policy strategy stabilizes policy entropy, almost completely suppressing its decline. Although it maintains strong exploration capability, it simultaneously impedes effective policy convergence, resulting in limited performance gains. The Off-policy + Entropy Loss strategy suffers from training instability issues. While this method maintains a gradual entropy decline in early training, as training progresses and policy entropy drops below a certain threshold, the magnitude of entropy regularization loss gradually increases and surpasses the GRPO main loss to become the dominant term, triggering abnormal policy entropy increases that ultimately lead to training collapse. Furthermore, this method requires extensive hyperparameter tuning experiments to configure the entropy loss weighting coefficient, increasing application costs. In summary, the Clip-higher strategy achieves an optimal balance between exploration capability and convergence efficiency by moderately controlling the decay rate of policy entropy, making it the best choice for the scenario in this study.
Importance Sampling Granularity Exploration. The standard GRPO algorithm computes importance sampling ratios at the token level, i.e., independently calculating for each position t as shown in equation 14. Inspired by the ”sentence-level reward” approach in GSPO(GSPO) and GMPO(GMPO), we explored sentence-level importance sampling, which computes a unified importance weight for the entire output sequence based on sequence likelihood as shown in equation 21. This weight is shared by all tokens in the sequence, ensuring that the entire reasoning chain is consistently reinforced or suppressed as a whole.
@@ 3.5.2 Benchmark Construction
This subset was designed to provide challenging data to differentiate the capabilities of advanced LLMs. We focused on mining hard cases, which we categorize into explicit demands (Multi-Attribute, MR) and implicit demands. The implicit demands include four distinct types: Negation (NE), Alternatives (AL), Knowledge-dependent (KD), and Reasoning-dependent (RD). The mining process was tailored for these types. NE and AL queries were filtered using predefined keywords and regular expressions. KD and RD queries, which often appear as long-tail queries, were mined by first undersampling low-frequency logs and then using a powerful LLM (Qwen3-235B) with a self-consistency filter (requiring ¿4 identical results across 8 inferences). MR queries were identified using an NER model on length-filtered queries to find those with more than 5 attributes. To ensure the final set was genuinely difficult, all candidates were filtered through a baseline relevance model; we retained only those samples where the model’s prediction was correct five or fewer times across eight inferences, indicating high model uncertainty.
@@ 4 Experiment
In this section, we conduct a systematic validation and analysis of the proposed methods on the RAIR benchmark. Our experiments are divided into two parts: first, through quantitative metrics on benchmarks, we validate the following two core questions: (1) the magnitude of performance improvement brought to the model by fine tuning; (2) the performance advantages demonstrated by the proposed LORE model across different data distribution subsets, including general samples, hard samples, and visually salient samples. Second, we employ case studies to illustrate the new capabilities that our methods confer upon the model and the types of problems they enable it to solve.
@@ 4.1.2 Baseline
As comparison baselines, we select the following two categories of models: (1) Top-tier large language models (open-source and closed-source): All models are evaluated using zero-shot direct prompting with a unified prompt template. The prompts comprise the complete relevance task definition and discrimination rules. The inference configuration follows the best practices reported in the official documentation of each model: employing single-pass inference with temperature parameters set to the officially recommended optimal values for each model (e.g., 0.7 for GPT-4, 0.6 for Llama-3, etc.), ensuring that each model is compared under its optimal operating conditions. (2) Vanilla SFT Model: This model is trained using only labels for SFT, without incorporating CoT modeling or the RL phase, serving to validate the joint contribution of CoT modeling and reinforcement learning optimization in our method.
@@ 4.2 Main Result
Qwen2.5-7B(base)(Qwen3) 0.775 0.689 0.391 0.531 0.382 0.312
@@ 5 Application
(2) Knowledge Distillation: for medium-frequency query-item pairs (approximately 65% of total traffic), which constitute a substantial proportion. Pre-computing and caching all query-item discrimination results is neither feasible nor cost-effective. Therefore, we maintained the original ”coarse ranking-fine ranking” processing pipeline. Meanwhile, the LLM was leveraged to generate large-scale, high-quality training data, systematically enhancing the discrimination capabilities of both online coarse ranking and fine ranking models through knowledge distillation techniques.through knowledge distillation techniques.
@@ 6 Discussion
During CoT data distillation experiments, we observed that distilling synthetic CoT data paradoxically degrades model performance compared to vanilla SFT. As show in Table 12, the cold-start model exhibits a 4.2% decline in pass@1 metrics relative to Naive-SFT. This degradation stems from training-inference distribution shift: the model relies on high-quality contexts generated by the teacher during training, whereas inference utilizes its own autoregressive outputs as subsequent-step inputs. Such discrepancy triggers error accumulation and propagation through the reasoning chain, ultimately causing misjudgments. Notably, CoT-distilled models demonstrate significant improvements in pass@8 metrics (2.7% increase), indicating that distillation enhances the upper bound of model capability by expanding answer space diversity and injecting multi-step reasoning proficiency.
@@
Studies such as DeepSeek-R1 have observed the emergent phenomenon of long chains-of-thought and the model’s ”Aha Moment.” We similarly conducted a tracking analysis of the dynamic changes in model output length during the reinforcement learning process, with results presented in Fig 17. The experimental findings reveal that as model capability continues to improve, the output length exhibits an overall decreasing trend and eventually stabilizes. This phenomenon suggests that long chains-of-thought are not a necessary condition for model capability enhancement, but rather an accompanying phenomenon during the training process. Further analysis indicates that different task types exhibit significant variations in their requirements for reasoning chain length: tasks such as mathematical computation and code generation typically necessitate longer step-by-step reasoning processes. For the relevance task, accurately identifying query intent and item attributes, and reasoning based on discrimination rules, constitutes an intuitive and clear chain of thought. Redundant reasoning processes beyond this do not yield performance improvements.

### v2

@@ 3.3.3 Distillation by SFT
Fig 9 illustrates the trends in model metrics as the data proportion increases. The results demonstrate that the model can rapidly acquire the correct reasoning format in the early training stage, achieving a format accuracy exceeding 98% with only 20% of the training data. The Pass@8 metric exhibits rapid growth in the early training phase but subsequently displays a phenomenon of diminishing marginal returns. When approximately 60% of the training data is utilized, the growth rate of this metric plateaus, showing no significant improvement. Based on these experimental observations, this study selects the 60% data proportion as the optimal training scale, with which the cold-start model is trained.
@@ 3.4.3 Training Strategy Optimization
Dynamic Sampling.(DAPO) As RL training progresses, sample effectiveness gradually decays: after the model policy converges, an increasing number of samples exhibit consistent outcomes across multiple sampling attempts (either all correct or all incorrect), resulting in an advantage function $A_{i}=0$. Although these saturated samples participate in training, they fail to generate effective gradients, leading to computational resource waste and slowing the convergence speed. To address this issue, we introduce a dynamic sampling strategy: during the rollout phase, we calculate sample advantage values in real-time and discard zero-advantage samples. A policy update round is only triggered when the accumulated number of non-zero advantage samples reaches the preset batch size.
Entropy Collapse Optimization. During RL training, as the policy gradually converges, the entropy of the model’s output distribution rapidly decreases, causing generation results to become deterministic. This premature stabilization hinders effective exploration of the reasoning manifold, making it difficult to discover better reasoning paths and ultimately constraining the improvement of the performance upper bound. To address this issue, we explored multiple strategies to slow down the entropy decline and maintain the model’s exploration capability.
@@
(2) On-policy. Traditional off-policy methods perform multi-epoch training on the same batch of experience data, using importance sampling constraints to prevent excessive policy deviation. However, this conservative update mechanism limits the model’s exploration capability. To address this, we adopt a strictly on-policy strategy: each batch of sampled data (batch_sizegroup_sizesample_numbatch\_size*group\_size*sample\_num) is discarded after being used for only one gradient update, and the next update uses freshly sampled data from the new policy.
@@
$\begin{split}\mathcal{J}_{\text{GRPO}}^{2}(\theta)=&\mathbb{E}_{x\sim\mathcal{D},\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|x)}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|y_{i}|}\sum_{t=1}^{|y_{i}|}\min\left(w_{i,t}(\theta)\widehat{A}_{i,t},\text{clip}(w_{i,t}(\theta),1-\varepsilon,1+\varepsilon)\widehat{A}_{i,t}\right)\right]\\ &\quad+\alpha\cdot\mathbb{E}_{x\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta}(\cdot|x)}\left[-\frac{1}{G}\sum_{i=1}^{G}\log\pi_{\theta}(o_{i}|x)\right]\end{split}$ (20)
@@
Fig 13 presents a comparison of several methods. The Clip-higher strategy exhibits optimal performance by effectively slowing the decline rate of policy entropy while maintaining its decreasing trend, achieving a favorable balance between exploration and exploitation, and ensuring training stability alongside significant performance improvements. The On-policy strategy stabilizes policy entropy, almost completely suppressing its decline. Although it maintains strong exploration capability, it simultaneously impedes effective policy convergence, resulting in limited performance gains. The off-policy + entropy loss strategy suffers from training instability issues. Although this method maintains a gradual entropy decline in early training, as training progresses and policy entropy drops below a certain threshold, the magnitude of entropy regularization loss gradually increases and surpasses the GRPO main loss to become the dominant term, triggering abnormal policy entropy increases that ultimately lead to training collapse. Furthermore, this method requires extensive hyperparameter tuning experiments to configure the entropy loss weighting coefficient, increasing application costs. In summary, the Clip-higher strategy achieves an optimal balance between exploration capability and convergence efficiency by moderately controlling the decay rate of policy entropy, making it the best choice for the scenario in this study.
Importance Sampling Granularity Exploration. The standard GRPO algorithm computes importance sampling ratios at the token level, i.e., independently calculating for each position t as shown in equation 14. Inspired by GSPO(GSPO) and GMPO(GMPO), we explored sentence-level importance sampling, which computes a unified importance weight for the entire output sequence based on sequence likelihood as shown in equation 21. This weight is shared by all tokens in the sequence, ensuring that the entire reasoning chain is consistently reinforced or suppressed as a whole.
@@ 3.5.2 Benchmark Construction
This subset was designed to provide challenging data to differentiate the capabilities of advanced LLMs. We focused on mining hard cases, which we categorize into explicit demands (Multi-Attribute, MA) and implicit demands. The implicit demands include four distinct types: Negation (NE), Alternatives (AL), Knowledge-dependent (KD), and Reasoning-dependent (RD). The mining process was tailored for these types. NE and AL queries were filtered using predefined keywords and regular expressions. KD and RD queries, which often appear as long-tail queries, were mined by first undersampling low-frequency logs and then using a powerful LLM (Qwen3-235B) with a self-consistency filter (requiring >> 4 identical results across 8 inferences). MR queries were identified using an NER model on length-filtered queries to find those with more than 5 attributes. To ensure the final set was genuinely difficult, all candidates were filtered through a baseline relevance model; we retained only those samples where the model’s prediction was correct five or fewer times across eight inferences, indicating high model uncertainty.
@@ 4 Experiment
In this section, we conduct a systematic validation and analysis of the proposed methods on the RAIR benchmark. Our experiments are divided into two parts: first, through quantitative metrics on benchmarks, we validate the following two core questions: (1) the magnitude of performance improvement brought to the model by fine tuning; (2) the performance advantages demonstrated by the proposed LORE model across different data distribution, including general samples, hard samples, and visually salient samples. Second, we employ case studies to illustrate the new capabilities that our methods confer upon the model and the types of problems they enable it to solve.
@@ 4.1.2 Baseline
As comparison baselines, we select the following two categories of models: (1) Top-tier large language models (open-source and closed-source): All models are evaluated using zero-shot direct prompting with a unified prompt template. The prompts comprise the complete relevance task definition and discrimination rules. The inference configuration follows the best practices reported in the official documentation of each model: employing single-pass inference with temperature parameters set to the officially recommended optimal values for each model (e.g., 0.7 for GPT-4, 0.6 for Llama-3, etc.), ensuring that each model is compared under its optimal operating conditions. (2) Vanilla SFT Model: This model is trained using only labels, without incorporating CoT modeling or the RL phase, serving to validate the joint contribution of CoT modeling and reinforcement learning optimization in our method.
@@ 4.2 Main Result
Qwen2.5-7B(base)(qwen25) 0.775 0.689 0.391 0.531 0.382 0.312
@@ 5 Application
(2) Knowledge Distillation: for medium-frequency query-item pairs (approximately 65% of total traffic), which constitute a substantial proportion. Pre-computing and caching all query-item discrimination results is neither feasible nor cost-effective. Therefore, we maintained the original ”coarse ranking-fine ranking” processing pipeline. Meanwhile, the LLM was leveraged to generate large-scale, high-quality training data, systematically enhancing the discrimination capabilities of both online coarse ranking and fine ranking models through knowledge distillation techniques.through knowledge distillation techniques(ELLM).
@@ 6 Discussion
In our CoT data distillation experiments, we observed that distilling synthetic CoT data paradoxically degrades model performance compared to vanilla SFT. As shown in Table 12, the distilled model exhibits a 4.2% decline in pass@1 metrics relative to the Naive-SFT baseline. This degradation stems from the training-inference distribution shift: during training, the model conditions on high-quality contexts generated by the teacher, whereas during inference, it relies on its own autoregressive outputs. This discrepancy triggers error accumulation through the reasoning chain, ultimately leading to incorrect predictions. Notably, however, the distilled models demonstrate significant improvements in pass@8 metrics (2.7% increase), indicating that distillation enhances the model’s capability upper bound by expanding answer space diversity and instilling multi-step reasoning proficiency.
@@
Studies such as DeepSeek-R1 have observed the emergent phenomenon of long CoT and the model’s ”Aha Moment.” We similarly conducted a tracking analysis of the dynamic changes in model output length during the reinforcement learning process, with results presented in Fig 17. The experimental findings reveal that as model capability continues to improve, the output length exhibits an overall decreasing trend and eventually stabilizes. This phenomenon suggests that long CoT are not a necessary condition for model capability enhancement, but rather an accompanying phenomenon during the training process. Further analysis indicates that different task types exhibit significant variations in their requirements for reasoning chain length: tasks such as mathematical computation and code generation typically necessitate longer step-by-step reasoning processes. For the relevance task, accurately identifying query intent and item attributes, and reasoning based on discrimination rules, constitutes an intuitive and clear chain of thought. Redundant reasoning processes beyond this do not yield performance improvements.
@@ 9 Contributors
Core Contributors
Algorithm: Chenji Lu, Zhuo Chen, Hui Zhao, Zhiyuan Zeng, Ximing Zhang, Gang Zhao, Junjie Ren, Ruicong Xu, Haoran Li, Songyan Liu, Zhenyi Wang, Pengjie Wang, Jian Xu, Bo Zheng
Engineering: Qingtao Zeng, Huimin Yi, Siqi Luo, Kaiwen Wei, Chen Chen, Fei Wang, Xiaojie Zhang, Qifeng Li, Daisong Wang, Qing Xue, Jiawen Liao, Zexin Yan, Zhenyuan Lai, Peng Sun, Jiangwei Yu, Zhige Chen, Yang Wang, Rong Huang, Shihao Chen, Xiang Gao, Ju Huang, Weixun Wang, Zhendong Li, Feilei Du, Jiamang Wang
Annotation: Weihao Guo, Junjing Su, Nan Lin, Xinyao Zhou, Annotation Team